# Modeling Quantum Noise in Nanolasers using Markov Chains

### v1

@@ V Numerical comparison
In the previous section, we derived the LMC and showed how it converges to the typical laser rate equations with Langevin forces in the many-photon limit. In this section, we compare the numerical solutions to the LMC obtained through Gillespie’s method and the tau-leaping method with the numerical solution to the stochastic differential equations constituted by the laser rate equations with stochastic Langevin forces (see appendix C for numerical details as well as Ref. git for an open-source implementation). For reference, we also include the analytical small-signal analysis of the Langevin equations, which, in general, apply when the variations in photon and carrier number constitute small relative changes to the steady-state values (see appendix D for details).
@@ V.1 Small number of emitters ($n_{0}=1$)
In contrast, both the analytical and numerical solutions to the Langevin equation deviate from the ME. The deviation of the analytical solution is less severe and mostly concerns the $g^{(2)}(0)$-function below threshold and the photon number around lasing. The former is a well-known shortcoming of the Langevin approach Bundgaard-Nielsen et al. (2023), while the latter is more subtle. As discussed and shown in Appendix E, the deviation originates from the assumption $\expectationvalue{n_{e}n_{p}}=\expectationvalue{n_{p}}\expectationvalue{n_{e}}$ made in the LRE, which is generally not the case. This anti-correlation between the emitter and photon number leads to a difference in the mean-photon number, which is perfectly captured by the LMC model.
For the numerical solution of the Langevin approach, the deviations are more severe, and it is basically invalid for all pump powers, even for the mean photon number. One of the causes for this large deviation between the analytical and numerical solution to the Langevin equations is the possibility of stochastic forces leading to negative emitter or photon populations. To prevent this and to ensure numerical stability, any stochastic force that tries to reduce a population below zero is instead clamped to zero. The effect of this is a skewing of the noise profile, which causes inaccurate mean populations and statistics. In chemical modeling, this is a known issue and renders the numerical solution to the LRE invalid Anderson et al. (2019). This problem, of course, only occurs when the populations are very small, but since this is always the case below the lasing threshold, we see that, independently of the emitter numbers $n_{0}$, the numerical solutions are inaccurate below threshold. Specifically for $n_{0}=1$, the population is so small for all pump rates that for no pump value, the numerical solution to the LRE is accurate. In appendix C, we explore other ways of handling negative populations, but none of these solve the problem.
@@ V.2 Intermediate number of emitters ($n_{0}=10$)
In Fig. 2, for $n_{0}=10$, it is computationally infeasible to naively perform calculations using the ME approach. For identical emitters, one can, however, slightly extend the simulation regime by utilizing the symmetry of the identical emitters. Using the so-called Permutational Invariant Quantum Solver (PIQS) Shammah et al. (2018) implemented in Qutip Lambert et al. (2024), we are able to simulate almost the entirety of this regime, with the exception of when the photon number is too large. As we will see in the following section, the numerical cost of the ME is, however, large: more than two orders of magnitude larger than the Gillespie approach.
As we increase the emitter number to $n_{0}=10$, we observe lasing although the threshold transition is not sharp and quenching effects terminate the lasing regime at higher pump rates Mu and Savage (1992); Bundgaard-Nielsen et al. (2023, 2025). Gillespie’s method, tau-leaping, and analytical Langevin approaches show excellent agreement with the ME results. We observe only a minor deviation in the analytical Langevin $g^{(2)}(0)$ function near the lasing threshold, likely due to the breakdown of the mean-field approximation $\expectationvalue{n_{e}n_{p}}\neq\expectationvalue{n_{p}}\expectationvalue{n_{e}}$ Bundgaard-Nielsen et al. (2023) as also discussed in the last section. The numerical solution of the Langevin equations is still invalid for all pump powers for the same reason as above.
@@ V.3 Large number of emitters ($n_{0}=100$10001000)
In Fig. 2, for $n_{0}=100$, we are approaching a larger laser system, which is impossible to simulate using a ME, even when exploiting symmetries. For larger emitter and photon numbers, we thus do not have ME results, and instead establish the validity of our comparisons through two means. First, both previous work and the present study demonstrate excellent agreement between Gillespie’s method and ME calculations for systems with up to seven to ten emitters Bundgaard-Nielsen et al. (2025). This provides confidence in the accuracy of Gillespie’s method within this parameter range, and suggests that it would also be accurate for larger emitter numbers. Second, we focus on the relative performance and systematic differences between methods, which remain meaningful even without absolute ground truth verification. These comparisons reveal regimes where each approach excels or fails, providing useful insights.
With $n_{0}=100$ and $\beta=0.01$, the transition to lasing is now marked by a sharp kink in the mean photon number as a function of pump rate, as well as the sudden transition of the $g^{(2)}(0)$-function from 2 to 1. We now see that the numerical Langevin approach agrees with the analytical Langevin approach above threshold, which is due to the large photon population obtained above threshold. The numerical solution below the threshold is still invalid due to the low photon number. The tau-leaping and analytical Langevin approach agree very well for all quantities at all pump rates. Gillespie’s method shows good agreement except in the RIN above the threshold. This deviation is due to an insufficient number of samples, which limits the statistical accuracy. Due to practical reasons, we have limited the number of steps taken in the Markov chain to $2\cdot 10^{10}$, which in this case turns out to be insufficient to accurately sample the photon statistics. This highlights a different regime where the tau-leaping approach is more efficient than Gillespie’s approach, since the tau-leaping approach permits multiple events to occur at each timestep as long as they do not perturb the population significantly, i.e., the populations of emitters and photons are large enough that the changes occurring during the time step do not significantly change the rates appearing in Table 2. For large photon populations, this means that within a tau-leap, a significant number of events happen, which allows the tau-leaping to sample much more efficiently than Gillespie’s approach, which only allows a single event to occur per step. Previously, it was thought that Gillespie’s method is always more efficient than the tau-leaping approach André et al. (2020), but we here show that it depends on the parameter regime. We highlight this further in section VI, where we compare the computational efficiency of the different numerical approaches.
@@
Figure 3: Overview of the validity of methods to compute laser quantum noise as a function of the number of photons and emitters in the laser. The boxes of different colors indicate the areas in which the different methods are valid or computationally feasible. The crossed-out area marks a combination of photon and emitter numbers, which with the parameters used in this paper, were not obtainable.
@@ V.4 Performance Summary
We define the relative error as $\delta^{2}=\max({\expectationvalue{\delta_{g^{(2)}(0)}^{2}}},{\expectationvalue{\delta_{RIN}^{2}}},{\expectationvalue{\delta_{n_{p}}^{2}}})$, where $\delta_{X}=X_{sim}-X_{truth}$ is the difference between the simulation results $X_{sim}$ (of 5 independent runs) and the ground truth value $X_{truth}$, which for $n_{0}=1$ and $n_{0}=10$ is assumed to be the ME results, and for $n_{0}=100$ and $n_{0}=10000$ taken to be the analytical solution to the langevin approaches. We vary the wall-clock time (the amount of time the simulation has run in real time) by adjusting the number of numerical steps taken in the simulation. The more numerical steps, means more samples and thus the more accurate averages, which leads to a lower relative error.
@@
For $n_{0}=100$, we have picked parameters where we are above threshold with many photons, and thus the numerical Langevin approach is accurate and is also seen to be more efficient in computational time, when compared with the other approaches. As discussed in the previous section, Gillespie’s approach starts to be inefficient as it is simply taking too small steps by only allowing a single event to happen. This is clearly seen in the high relative error, even for long computational times. At the same time, Gillespie’s approach is, in principle, accurate, which is seen in the steady increase in accuracy as the computational time increases. The tau-leaping approach, on the other hand, is more efficient and has a lower relative error, since it allows multiple events to occur within each step. It is still more computationally intensive when compared with the numerical Langevin approach, however, which motivates the formulation of stochastic differential equations: they are quicker to solve. For $n_{0}=10000$, we see a similar result, although now the numerical Langevin approach is just much more efficient than other approaches, and even the tau-leaping approach struggles.
We note that we have picked only particular pump values at each emitter number value. Thus, the conclusion that for $n_{0}>100$, the LRE outperforms the tau-leaping approach is not in general true. Specifically, as mentioned previously, if we go below the threshold, the numerical solution to the LRE is invalid, no matter the number of steps, and the appropriate approach would thus here be the tau-leaping. Still, this section highlights the general trend that for larger laser systems, the LRE becomes more and more efficient when compared with the tau-leaping and Gillespie approach. However, to have an accurate modeling of the complete transition from below to above lasing it is necessary to use the Gillespie or tau-leaping approaches.
@@ VII Outlook and Conclusion
In this paper, we have established the formal connection between the ME and the so-called stochastic LMC, which treats laser populations as discrete integer variables. We furthermore showed that in the many-photon limit, the LMC simplifies to the LRE, thereby establishing the validity regimes of different approaches to quantum noise in nanolasers.
Our numerical comparison of accuracy and computational efficiency revealed that each method has distinct regimes of applicability. Gillespie’s method provides the most efficient sampling of the LMC for small populations, while tau-leaping becomes superior for intermediate system sizes. The numerical Langevin approach is only accurate and efficient for very large populations where quantum discreteness effects become negligible.
We also showcase one of the main problems with solving the LRE numerically; below threshold, when the photon population is small, the Gaussian white noise can cause negative populations. This has to be dealt with by limiting or preventing the solver from producing negative populations, which ultimately skews the noise distribution and causes inaccurate statistics. The analytical solution to the LRE works for the macroscopic regimes considered in this paper. However, if the laser rate equations are more complicated or large signal modulation needs to be studied, analytical solutions can be difficult, if not impossible, to obtain. The LMC method, on the other hand, has no such restrictions and is numerically stable: it can be used to explore more complex laser systems involving, e.g., non-linear mirrors Sloan et al. (2024); Nguyen et al. (2023); Rivera et al. (2023); Pontula et al. (2022) or nanolaser modulation, both interesting research directions.
@@ Appendix C Numerical details
For the simulations of Fig. 2, a number of smaller numerical details need examination: how to choose the time step in the tau-leaping and numerical solution to the LRE, how averages are obtained, how errors are estimated, and how negative populations are handled in the numerical solution to the LRE. In this appendix, we address the above. We note that there is also an open-source implementation available to produce all the results of the paper in: git .

### v2

@@ V Numerical comparison
In the previous section, we derived the LMC and showed how it converges to the typical laser rate equations with Langevin forces in the many-photon limit. In this section, we compare the numerical solutions to the LMC obtained through Gillespie’s method and the tau-leaping method with the numerical solution to the stochastic differential equations constituted by the laser rate equations with stochastic Langevin forces (see appendix C for numerical details as well as Ref. git (2026) for an open-source implementation). For reference, we also include the analytical small-signal analysis of the Langevin equations, which, in general, apply when the variations in photon and carrier number constitute small relative changes to the steady-state values (see appendix D for details).
@@ V.1 Small number of emitters ($n_{0}=1$)
In contrast, both the analytical and numerical solutions to the Langevin equation deviate from the ME. The deviation of the analytical solution is less severe and mostly concerns the $g^{(2)}(0)$-function below threshold and the photon number around lasing. The former is a well-known shortcoming of the Langevin approach Bundgaard-Nielsen et al. (2023), while the latter is more subtle. As discussed and shown in Appendix E, the deviation originates from the mean-field approximation $\expectationvalue{n_{e}n_{p}}=\expectationvalue{n_{p}}\expectationvalue{n_{e}}$ made in the LRE, which is generally not the case. This anti-correlation between the emitter and photon number leads to a difference in the mean-photon number, which is perfectly captured by the LMC model.
For the numerical solution of the Langevin approach, the deviations are more severe, and it is basically invalid for all pump powers, even for the mean photon number. One of the causes for this large deviation between the analytical and numerical solution to the Langevin equations is the possibility of stochastic fluctuations leading to negative emitter or photon populations. To prevent this and to ensure numerical stability, any stochastic fluctuation that tries to reduce a population below zero is instead clamped to zero. The effect of this is a skewing of the noise profile, which causes inaccurate mean populations and statistics. In chemical modeling, this is a known issue and renders the numerical solution to the LRE invalid Anderson et al. (2019). This problem, of course, only occurs when the populations are very small, but since this is always the case below the lasing threshold, we see that, independently of the emitter numbers $n_{0}$, the numerical solutions are inaccurate below threshold. Specifically, for $n_{0}=1$, the population is so small for all pump rates that for no pump value, the numerical solution to the LRE is accurate. In appendix C, we explore other ways of handling negative populations, but none of these solve the problem.
@@ V.2 Intermediate number of emitters ($n_{0}=10$)
As we increase the emitter number to $n_{0}=10$ in Fig. 2, we observe lasing although the threshold transition is not sharp and quenching effects terminate the lasing regime at higher pump rates Mu and Savage (1992); Bundgaard-Nielsen et al. (2023, 2025).
Here, for $n_{0}=10$, it is computationally unfeasible to perform calculations using the standard ME approach. For identical emitters, one can, however, slightly extend the simulation regime by utilizing the symmetry of the identical emitters. Using the so-called Permutational Invariant Quantum Solver (PIQS) Shammah et al. (2018) implemented in Qutip Lambert et al. (2024), we are able to simulate most of the regime, except when the photon number is too large. As we will see in the following section, the numerical cost of the ME is, however, large: simulation times are more than two orders of magnitude larger than the Gillespie approach.
Gillespie’s method, tau-leaping, and analytical Langevin approaches show excellent agreement with the ME results. We observe only a minor deviation in the analytical Langevin $g^{(2)}(0)$ function near the lasing threshold, likely due to the breakdown of the mean-field approximation $\expectationvalue{n_{e}n_{p}}\neq\expectationvalue{n_{p}}\expectationvalue{n_{e}}$ Bundgaard-Nielsen et al. (2023) as also discussed in the last section. The numerical solution of the Langevin equations is still invalid for all pump powers for the same reason as above.
@@ V.3 Large number of emitters ($n_{0}=100$10001000)
In Fig. 2, for $n_{0}=100$, we are approaching a larger laser system, which is impossible to simulate using a ME, even when exploiting symmetries. For larger emitter and photon numbers, we thus do not have ME results, and instead establish the validity of our comparisons by two means. First, both previous work and the present study demonstrate excellent agreement between Gillespie’s method and ME calculations for systems with up to seven to ten emitters Bundgaard-Nielsen et al. (2025). This provides confidence in the accuracy of Gillespie’s method within this parameter range, and suggests that it would also be accurate for larger emitter numbers. Second, we focus on the relative performance and systematic differences between methods, which remain meaningful even without absolute ground truth verification. These comparisons reveal regimes where each approach excels or fails, providing useful insights.
With $n_{0}=100$ and $\beta=0.01$, the transition to lasing is now marked by a sharp kink in the mean photon number as a function of pump rate, as well as the sudden transition of the $g^{(2)}(0)$-function from 2 to 1. We now see that the numerical Langevin approach agrees with the analytical Langevin approach above threshold, which is due to the large photon population obtained above threshold. The numerical solution below the threshold is still invalid due to the low photon number. The tau-leaping and analytical Langevin approach agree very well for all quantities at all pump rates. Gillespie’s method shows good agreement except in the RIN above the threshold. This deviation is due to insufficient statistical sampling. For practical reasons, we limit the Markov-chain simulations to $2\times 10^{10}$ steps. In the Gillespie approach, large photon numbers lead to very short waiting times between individual events, so that even this number of steps corresponds to a limited simulation time and therefore does not adequately sample the photon statistics. This highlights a different regime where the tau-leaping approach is more efficient than Gillespie’s approach, since the tau-leaping approach permits multiple events to occur at each timestep as long as they do not perturb the population significantly, i.e., the populations of emitters and photons are large enough that the changes occurring during the time step do not significantly change the rates appearing in Table 2. For large photon populations, this means that within a tau-leap, a significant number of events happen, which allows the tau-leaping to sample much more efficiently than Gillespie’s approach, which only allows a single event to occur per step. Previously, it was thought that Gillespie’s method is always more efficient than the tau-leaping approach André et al. (2020), but we here show that it depends on the parameter regime. We highlight this further in section VI, where we compare the computational efficiency of the different numerical approaches.
@@
Figure 3: Overview of the validity of methods to compute laser quantum noise as a function of the number of photons and emitters in the laser. The boxes of different colors indicate the areas in which the different methods are valid or computationally feasible. The crossed-out area marks a combination of photon and emitter numbers that are not accessible with the parameters used in this paper due to the physical effect of pump-induced quenching.
@@ V.4 Performance Summary
We define the relative error as $\delta^{2}=\max({\expectationvalue{\delta_{g^{(2)}(0)}^{2}}},{\expectationvalue{\delta_{RIN}^{2}}},{\expectationvalue{\delta_{n_{p}}^{2}}})$, where $\delta_{X}=X_{sim}-X_{truth}$ is the difference between the simulation results $X_{sim}$ (of 5 independent runs) and the ground truth value $X_{truth}$, which for $n_{0}=1$ and $n_{0}=10$ is assumed to be the ME results, and for $n_{0}=100$ and $n_{0}=10000$ taken to be the analytical small-signal analysis of the LRE. We vary the wall-clock time (the amount of time the simulation has run in real time) by adjusting the number of numerical steps taken in the simulation. The more numerical steps, means more samples and thus the more accurate averages, which leads to a lower relative error.
@@
For $n_{0}=100$, we have picked parameters where we are above threshold with many photons, and thus the numerical Langevin approach is accurate and is also seen to be more efficient with respect to computational time, when compared with the other approaches. As discussed in the previous section, Gillespie’s approach starts to be inefficient as it is simply taking too small steps by only allowing photon and emitter numbers to change by one unit during a time-step. This is clearly seen in the high relative error, even for long computational times. At the same time, Gillespie’s approach is, in principle, accurate, which is seen in the steady increase in accuracy as the computational time increases. The tau-leaping approach, on the other hand, is more efficient and has a lower relative error, since it allows multiple events to occur within each step. It is still more computationally intensive when compared with the numerical Langevin approach, however, which motivates the formulation of stochastic differential equations: they are quicker to solve. For $n_{0}=10000$, we see a similar result, although now the numerical Langevin approach is just much more efficient than other approaches, and even the tau-leaping approach struggles.
We note that we have picked only particular pump values at each emitter number value. Thus, the conclusion that for $n_{0}>100$, the LRE outperforms the tau-leaping approach is not in generally valid. Specifically, as mentioned previously, if we go below the threshold, the numerical solution to the LRE is invalid, no matter the number of steps, and the appropriate approach would thus here be the tau-leaping. Still, this section highlights the general trend that for larger laser systems, the LRE becomes more and more efficient when compared with the tau-leaping and Gillespie approach. However, to have an accurate modeling of the complete transition from below to above lasing it is necessary to use the Gillespie or tau-leaping approaches.
@@ VII Outlook and Conclusion
In this paper, we have established the formal connection between the master equation and the so-called stochastic Laser Markov Chain, which treats laser populations as discrete integer variables. We furthermore showed that in the many-photon limit, the Laser Markov Chain simplifies to the Langevin Rate Equations, thereby establishing the validity regimes of different approaches to quantum noise in nanolasers.
Our numerical comparison of accuracy and computational efficiency revealed that each method has distinct regimes of applicability. Gillespie’s method provides the most efficient sampling of the Laser Markov Chain for small populations, while tau-leaping becomes superior for intermediate system sizes. The numerical Langevin approach is only accurate and efficient for very large populations where quantum discreteness effects become negligible.
We also showcase one of the main problems with solving the Langevin Rate Equations numerically; below threshold, when the photon population is small, the Gaussian white noise can cause negative populations. This has to be dealt with by limiting or preventing the solver from producing negative populations, which ultimately skews the noise distribution and causes inaccurate statistics. The analytical solution to the Langevin Rate Equations works for the macroscopic regimes considered in this paper. However, if the laser rate equations are more complicated or large signal modulation needs to be studied, analytical solutions can be difficult, if not impossible, to obtain. The Laser Markov Chain method, on the other hand, has no such restrictions and is numerically stable: it can be used to explore more complex laser systems involving, e.g., non-linear mirrors Sloan et al. (2024); Nguyen et al. (2023); Rivera et al. (2023); Pontula et al. (2022) or nanolaser modulation, both interesting research directions. Another interesting direction is to include the emitter-photon polarization $\sigma^{\dagger}a$ directly in the Laser Markov Chain. Langevin equations, which include these correlations, have previously been derived Maki et al. (1989); Drummond and Raymer (1991), and thus it does not seem impossible to derive a Laser Markov Chain which includes such correlations.
@@ Appendix C Numerical details
For the simulations of Fig. 2, a number of smaller numerical details need examination: how to choose the time step in the tau-leaping and numerical solution to the LRE, how averages are obtained, how errors are estimated, and how negative populations are handled in the numerical solution to the LRE. In this appendix, we address the above. We note that there is also an open-source implementation available to produce all the results of the paper in: git (2026).